OpenStudy (anonymous):

Hi can someone help me to better understand Finding the slope of a line please?

OpenStudy (igreen):

Question?

OpenStudy (igreen):

Well, we can take any two points on the line and plug them into the slope formula. \(\sf m = \dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (igreen):

Or we can do \(\sf\dfrac{Rise}{Run}\)

OpenStudy (yanasidlinskiy):

For example, you have a question something like this: Find the equation of the straight line that has a slope m=4 and passes through the point (-1,-6) Do you know what our first step is? As has been stated above, the slope formula is \(y=mx+b\)

OpenStudy (igreen):

Eh, just look at my tutorial. http://openstudy.com/study#/updates/5457ae42e4b0a717ff61bd1d

OpenStudy (yanasidlinskiy):

Hahha, better than ever! "Eh, just look at my tutorial" LOL!

OpenStudy (yanasidlinskiy):

I meant slope intercept.

OpenStudy (anonymous):

Thank you! I wish I could medal both of you but it only allows 1. @YanaSidlinskiy I'll medal you on a different question :)

OpenStudy (yanasidlinskiy):

Oh, no. It's fine. As long as it helps. I honestly don't care about Medals:) Give it to who you wish:)